OpenStudy (anonymous):

what is the angle formed by 7:15 o'clock

OpenStudy (anonymous):

165 degrees

OpenStudy (anonymous):

how did you come up with 165

OpenStudy (anonymous):

there are 360 degrees in a circle and there are 12 hours on the face of a clock. Thus, the difference in degrees from each hour on the clock is 360/12 = 30 degrees. At 7 o'clock the hour hand is at 7 - meaning that it is 30*7=210 degrees displaced from the origin which I have set to be 12:00. and at 15 minutes into the hour the minute hand is at the 3. 210 - 3*30 = the angular distance between the hands = 210-90 = 120 degrees, sorry i made a mistake in my original solution. The distance between the hands of the clock at 7:15 is 120 degrees.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

so can i say that 7:15 is a obtuse angle or reflex angle

OpenStudy (anonymous):

you could say the 120 degrees between the hour hand and the minute hand to be obtuse, reflex implies an angle over 180 degrees. If you went the other way around from the hour to the minute hand you could have a reflex angle of 240 degrees.